OpenStudy (anonymous):

How do you solve this Find two consecutive integers such that the lesser added to three times the greater gives a sum of 46

OpenStudy (unklerhaukus):

the the smaller of the integres be \(n\) the other inter is consecutive so it is \(n+1\)

OpenStudy (unklerhaukus):

\[n+3(n+1)=46\]

OpenStudy (anonymous):

let x be the 1st integer(lesser) 3(x+1) be the second integer

OpenStudy (unklerhaukus):

solve for \(n\)

OpenStudy (anonymous):

How do you get two numbers then

OpenStudy (unklerhaukus):

\(n\) is the smaller of the integres the other is \(n+1\)

OpenStudy (unklerhaukus):

for some reason im not getting integer values though

OpenStudy (anonymous):

Yea I don't get it

OpenStudy (anonymous):

i guess the question is wrong

OpenStudy (anonymous):

The integers have to be even

OpenStudy (anonymous):

let the two consecutive numbers be n, n + 2 the question says if we add the lesser number \[\left( n \right)\] to three times the greater number \[3\left( n + 2 \right)\] we get 46 so in the form of equation we get \[\left[ \left( n \right) + 3\left( n +2 \right) \right] = 46\] now on solving it \[\left[ n + 3n +6 \right] = 46\] \[4n + 6 = 46\] \[4n = 40\] \[n = 10\] the numbers are 10 and 12

OpenStudy (anonymous):

Yea that's right thanks so much

OpenStudy (anonymous):

wheres my medal :D

OpenStudy (anonymous):

Oh sorry